Solve: 5a + 2=7
2x +3= 7
3b+ 6 = 12
4a + 1= 13

Answer by JBarnum(2146)

im guessing these are 4 different problems
5a+2=7 subtract 2
5a=5 divide by 5
a=1


2x+3=7 subtract 3
2x=4 divide by 2
x=2


3b+6=12 subtract 6
3b=6 divide by 3
b=2


4a+1=13 subtract 1
4a=12 divide by 4
a=3